he charges to Work in Process—Assembly Department for a period, together with information concerning production, are as follows. All direct materials are placed in process at the beginning of production. Work in Process—Assembly Department Bal., 9,000 units, 40% completed 72,360 To Finished Goods, 41,500 units ? Direct materials, 40,000 units @ $6.80 272,000 Direct labor 80,000 Factory overhead 40,450 Bal., ? units, 30% completed ? Cost per equivalent units of $6.80 for Direct Materials and $3.00 for Conversion Costs. a. Based on the above data, determine the different costs listed below. 1. Cost of beginning work in process inventory completed this period $ 2. Cost of units transferred to finished goods during the period $ 3. Cost of ending work in process inventory $ 4. Cost per unit of the completed beginning work in process inventory (Rounded to the nearest cent

managerial accounting
Answer #1

a)

Material Conversion cost Total
Cost of beginning work in process inventory completed this period 0 16200   16200

**Material is added at beginning of process thus is was 100% completed at beginning

**conversion Cost of beginning work in process inventory completed this period = cost per equivalent unit * % of beginning work in process completed this period

                    = 3 * (9000*60%)

                    = 3 *5400

                  = 16200

If 40% is completed at beginning then remaining 60% is completed during the period .

2)

Material Conversion cost Total
cost of beginning inventory 72360
Cost of units completed during the period 41500*6.8 = 282200 41500*3= 124500 406700
Total cost of units transferred to finished goods inventory 479060

3)

Material Conversion cost Total
cost of ending work in process inventory 7500*6.8= 51000 7500* 406700

4)

Cost per unit of the completed beginning work in process inventory

Beginning work in process 72360
Cost of beginning work in process completed this period (Part 1) 16200
Total cost 88560
Number of units 9000
Cost per unit $ 9.84 per unit